OpenStudy (anonymous):

If a cotton ball is dropped from 12 meters with air resistance, what will be the velocity and acceleration at t = 1.00 s?

OpenStudy (anonymous):

applying F.P.D (Newton's 2nd Law) (summation)Forces/cotton-ball = m.a (vectors over the Forces and a) where a is the acceleration W(vector) + R(vector) = m.a(vector) R is the air resistance since a cotton ball is massless then W = m.g = 0 x 10 = 0 N projection along direction of motion (vertically download) Rcos(pi) = m.a -R = m.a but again the cotton ball is massless then -R = 0 then R = 0 !! x = (1/2)a.t^2 + V(0)t + x(0) 12 = (1/2)a.(1)^2 + 0(1) + 0 a = 24 m/s^2 V = (integral) a dt = 24t + V(0) = 24t + 0 = 24t V(1) = 24 m/s^2 :p

OpenStudy (anonymous):

Thank you!